OpenStudy (anonymous):

These are Quadratic Equations. and the question asks us to solve the equation by finding the square roots. 3x squared = 15. that is all i understand.

OpenStudy (lgbasallote):

first step divide both sides by 3

OpenStudy (lgbasallote):

what do you get?

OpenStudy (anonymous):

1 and 5

OpenStudy (lgbasallote):

\[x^2 = 5\] now to get rid of the square in x just take square root of both sides \[\sqrt x^2 = \sqrt 5\] \[x = \sqrt 5\]

OpenStudy (lgbasallote):

do you get that?

OpenStudy (anonymous):

Yes! That was not as hard as I thought! Thank you very much!

OpenStudy (lgbasallote):

happy to help ^_^

OpenStudy (anonymous):

OK not another one. 36x squared = 256. so i divide each side by 36?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

note...dont put it in decimal form ;) if you dont get a whole number when you divide but put as fraction

OpenStudy (lgbasallote):

just put as fraction*

OpenStudy (anonymous):

1/7

OpenStudy (lgbasallote):

i dont thing 256/36 = 1/7

OpenStudy (lgbasallote):

i dont think*

OpenStudy (lgbasallote):

let me just show you how to do it \[36x^2 = 256\] \[x^2 = \frac{256}{36}\] both 256 and 36 are perfect squares so if i square root both sides... \[\sqrt x^2 = \sqrt{\frac{256}{36}}\] \[x = \frac{16}{6}\] 16 is the square root of 256 and 6 is the square root of 36

OpenStudy (anonymous):

OK thanks! This math is new to me!

OpenStudy (lgbasallote):

haha..it's cool :)) well manipulating it is cool lol